OpenStudy (anonymous):

Find all sollutions on [0,2pi): cosx=+/-1/sqrt(4)

OpenStudy (psymon):

So cos(x) = 1/2 or -1/2. Well, there are two values where cos(x) = 1/2 and 2 values where cos(x) = -1/2. Do you have much knowledge of the unit circle?

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

@zepdrix

OpenStudy (luigi0210):

At what values does cos=+/- 1/2?